OpenStudy (anonymous):

Fans & Medals !!!!! Which is an equation in point-slope form for the given point and slope? Point:(2, -6) ; Slope:-3/4

OpenStudy (anonymous):

point slope form \[(y - y_1) = m(x-x_1)\] (x1, y1) = point and m = slope

OpenStudy (campbell_st):

point slope form is \[y - y_{1} = m(x - x_{1})\] where \[m = slope~~~~point = (x_{1},~~y_{1})\]

OpenStudy (jojo4eva):

y=-3/4x-18/4 i believe

OpenStudy (jojo4eva):

are there any answer choices

OpenStudy (campbell_st):

well that looks like slope intercept form...@jojo4eva

OpenStudy (jojo4eva):

well then it would just be y+6=-3/4(x-2)